OpenStudy (brainzonly):

Very simple...and very tricky. XD 6 divided by 2(1+2)= ? @Kevin :D

OpenStudy (kevin):

6/ 2(1 +2) = 6/ 2(3) = 6/6 = 1

OpenStudy (kevin):

Go to other question :D

OpenStudy (brainzonly):

PEMDAS won't help you here. LOL I LOVE this one. NO!! WRONG LOL

OpenStudy (kevin):

Lol

OpenStudy (brainzonly):

XD!!!!!!!! Ok. to the point: Do multiplication and division LEFT TO RIGHT!!! So 6 divided by 2(1+2)= ? 6 divided by 2(3)=? 3(3) =? 9 LOL

OpenStudy (brainzonly):

I have a question: HOW IN THE HELL DID YOU GET INTO HIGH SCHOOL???????????????????

OpenStudy (kevin):

So it should be like this ?